OpenStudy (hannahc234):

Will fan and medal!!!!!!!! How many solutions does the system of equations have? 3x+12y=20 y=-1/4x+5/3 A. one B. two C. Infinitely many D. none

OpenStudy (hannahc234):

@jabez177

OpenStudy (shadowlegendx):

Solve both of them for y

OpenStudy (shadowlegendx):

aka you want to isolate it

OpenStudy (vickyisthesmartone123):

To clarify, I am rewriting your equations so it is clear: 3x + 12y = 20 y = -(1/4)x + (5/3) These equations both represent lines. And with these lines there are 3 possibilities: a. 0 solutions (lines are parallel, do not cross) b. 1 solution (lines cross at one x,y point) c. infinite solutions (lines are identical, laying on top of each other) ...but how do we figure out which it is? We examine them by setting them equal to each other to compare. First, we simplify the equations so that one of the variables is common in both. Since one equation is already "y=", we will rearrange the other to match: 3x + 12y = 20 can be changed to: (3/12)x + (12/12)y = (20/12) [whole equation divided by 12] (1/4)x + y = (5/3) [simplifying the fractions] y = -(1/4)x + (5/3) [moving (1/4)x to the other side] and as you can see, this matches the 2nd equation of y = -(1/4)x + (5/3). Since the two equations are identical, there are INFINITE SOLUTIONS. This is because you solve to 0=0 when making y=y: -(1/4)x + (5/3) = -(1/4)x + (5/3) -(1/4)x + (1/4)x = (5/3) - (5/3) 0=0 BUT even if they were not identical, you would still set them equal to each other and simplify as far as you could. AND if doing so eliminates all the variables but are left with just numbers (for example 0=3), you would have parallel lines with no solutions. OR if doing so solves to a variable solution (for example x=-3), then you have one solution and you would put the x solution into either equation to get your y. Hope that helps :)

OpenStudy (shadowlegendx):

Excellent answer @Vickyisthesmartone123

OpenStudy (hannahc234):

thank you so much!!!!!

OpenStudy (vickyisthesmartone123):

@ShadowLegendX Thank you! :) @HannahC234 Your welcome! :D